Calculate the number of Li+ ions in 6.24 g Li3PO4
.
1
Expert's answer
2020-10-16T08:57:37-0400

M(Li3PO4) = 116 g/mol

m(Li3PO4) = 6.24 g

"n = \frac{m}{M}"

"n(Li3PO4) = \frac{6.24}{116} = 0.053\; moles"

1 mole of Li3PO4 contains 3 moles of Li+ ions.

0.053 moles of Li3PO4 contains "3 \times 0.053 = 0.159" moles of Li+ ions.

Proportion:

1 mole of Li+ ions – "6.022 \times 10^{23} \;ions"

0.159 moles of Li+ ions – x

"x = 0.159 \times 6.022 \times 10^{23} ions = 0.957 \times 10^{23}\; ions"

Answer: 9.5×1022 ions